# Entanglement and correlations between local observables in de Sitter spacetime

### v1

@@ I Introduction
Beyond its implications for cosmology, this result emphasizes the need to distinguish sharply between entanglement and correlation in quantum field theory. The intuitive link between the two often stems from our experience with pure states, where all correlations between a subsystem and its complement are necessarily due to entanglement. In field theory, however, the situation is more subtle. In particular, the reduced state of any field mode that is localized in space—i.e., has compact spatial support—is necessarily mixed (this is a direct consequence of the Reeh-Schlieder theorem [6]. For mixed states, there is no straightforward relationship between correlation and entanglement.
The entanglement content of de Sitter space has been extensively explored in the literature [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 14, 26, 27, 28, 29, 30, 31, 32, 16, 33, 34, 16, 35, 31, 36, 37, 38, 39], raising the natural question of what new insights this article seeks to provide. Previous studies have approached the problem from diverse perspectives and using a variety of tools, each with its own strengths and limitations—and sometimes producing results that appear in tension with one another. One prominent approach involves quantifying the entanglement between a region and its complement using entanglement entropy [24]. This method has led to interesting findings, suggesting that the entire de Sitter space contains more entanglement than Minkowski space—a result we independently confirm using a different methodology.
@@
The analysis presented here uncovers structural aspects of entanglement in the cosmological patch of de Sitter spacetime that, to our knowledge, have not been previously explored. In particular, by employing the concept of partner modes [66, 67, 68, 69], we are able to characterize the subtle manner in which entanglement is distributed in the Bunch-Davies vacuum, revealing important differences from Minkowski spacetime. This, in turn, clarifies that the apparent tension between results from entanglement entropy and entanglement harvesting protocols is not a contradiction. On the contrary, we demonstrate that these findings are closely connected, and in fact, one is a consequence of the other.
For a summary of some of the results presented here, see [70]. The rest of the paper is organized as follows, In Sec. II, we summarize the tools that will be used in the rest of the paper. In Sec. IV, we analyze the von Neumann entropy of a single field degree of freedom. Sections V and VI analyze mutual information and entanglement, respectively, between pairs of compactly supported modes of the field. Finally, in Sec. VII we focus on the notion of partners: the system that codifies the correlations between a local mode and the rest of the field. We use this notion to uncover the distribution of entanglement between different local modes of a scalar field in the Bunch-Davies vacuum in Sec. VII.2. Finally, Sec. VIII summarizes our findings and discusses their implications for curved spacetime quantum field theory and observational signatures of entanglement in the early universe. Throughout this paper, we use units in which $\hbar=c=1$ and adopt the following convention for the Fourier transform $\tilde{h}(\vec{k})=\int d^{3}x\,h(\vec{x})e^{i\vec{k}\cdot\vec{x}}$.
@@ III The cosmological patch of de Sitter spacetime
The de Sitter group in four spacetime dimensions has ten independent Killing vector fields. Locally, all of them are isometries of PdS. However, because the Poincaré patch covers only a portion of de Sitter space, not all of these transformations correspond to global isometries of PdS. Only the subgroup that leaves the Poincaré patch invariant corresponds to global isometries of PdS (see, e.g., [4, Sec. IV C]). This subgroup forms a seven-dimensional Lie group, generated by three spatial translations, three rotations, and one additional isometry defined by the Killing vector field
@@
In the flat-space limit $H\to 0$ and $\mu\to 0$ (in this order), the terms $\mathcal{N}_{\Phi\Phi}$, $\mathcal{N}_{\Pi\Pi}$, and $\mathcal{N}_{\Phi\Pi}$ vanish, and (45)-(47) reduce to the smeared two-point functions of Minkowski spacetime, as given in Eq. (II.2.1).
@@
It is also worth noting that condition (50) is morally analogous to introducing an infrared cutoff $k_{c}$ on the modulus of the wave numbers in Fourier space. However, (50) is a somewhat milder constraint, as it corresponds to removing only the zero mode $\vec{k}=0$, rather than truncating all low-momentum modes $|\vec{k}|<k_{c}$. This remark serves to contextualize the rather drastic limitation on accessible degrees of freedom imposed by introducing an infrared cutoff.
@@ IV Von Neumann entropy
Case 1: Using Eqs. (45)- (47), we find $\nu_{I}^{2}-(\nu_{I}^{\mathrm{Mink}})^{2}=\mathfrak{a}\,(RH)^{4-4\mu^{2}}(1+\mathcal{O}(\mu^{2}))+\mathcal{O}[(RH)^{3-3\mu^{2}}]\,$, with
@@
Since homogeneous Sobolev spaces of order $|s|<3/2$ are Hilbert spaces with an inner product given in Eq. (92[82], the Cauchy-Schwarz inequality implies that $\mathfrak{a}\geq 0$. Saturation of this inequality, i.e., $\mathfrak{a}=0$, can only occur if $f_{A}^{(1)}=f_{A}^{(2)}$, which corresponds to Case 3, and will be analyzed below. Thus, $\nu_{A}^{2}-(\nu_{A}^{\mathrm{Mink}})^{2}>0$ for $RH\gg 1$, $mH\ll 1$ whenever $0\neq f^{(A)}_{I}\neq f^{(2)}_{A}\neq 0$. Furthermore, since $\mathfrak{a}>0$, $\nu_{A}^{2}-(\nu_{A}^{\mathrm{Mink}})^{2}$ —and hence $\nu_{A}$—grows with $H$, which automatically implies that the entropy of the mode $A$ increases as $H$ increases.
@@ V.1 Examples
Figure 7: Geometric configuration of the modes in Example 6. The shell has inner and outer radii $R_{S}\pm d$, where $d$ is its width and $R_{S}=R+\Delta x$. Here, $\Delta x$ denotes the radial distance between the edge of the ball (of radius $R$), and the inner edge of the shell.
@@ VI.1 Examples
This case is interesting because $\tilde{\nu}_{-}^{\mathrm{Mink}}<1$ for a certain range of $|\Delta\vec{x}|$ (see Fig. 10), meaning that the two modes are entangled in the Minkowski vacuum. In this ball-shell configuration, the two modes are “closer together” compared to the ball-ball configuration, since the shell completely surrounds the ball, which intuitively explains why there is more entanglement.
@@
    LN decays exponentially with the radial distance $|\Delta x|$ between the shell and ball modes, vanishing beyond a certain threshold distance;
@@ VII Partner systems and entanglement distribution
The apparent contradiction is resolved by analyzing how correlations and entanglement are spatially distributed in the Bunch–Davies vacuum. In particular, we show that the notion of a partner mode [66, 67, 68, 69] provides a convenient framework to characterize the spatial structure of entanglement in quantum field theory in the cosmological patch of de Sitter spacetime.
@@ VII.1 Asymptotic behavior of the partner of a local mode in de Bunch–Davies vacuum
More precisely, let $\gamma_{A_{p}}$ be any vector in $\Gamma_{A_{p}}^{\mathbb{C}}$—the complex version of $\Gamma_{A_{p}}$—with unit Klein–Gordon norm (see footnote 6), so that the pair $(\gamma_{A_{p}},\gamma_{A_{p}}^{*})$ forms a symplectically-orthonormal basis of $\Gamma^{\mathbb{C}}_{A_{p}}$, from which we can obtain a basis in $\Gamma_{A_{p}}$. If we write $\gamma_{A_{p}}(\vec{x})=(g_{p}(\vec{x}),f_{p}(\vec{x}))$, then $g_{p}(\vec{x})$ falls off at least as $r^{-2\mu^{2}}$ and $f_{p}(\vec{x})$ falls off as $r^{-1-\mu^{2}}$, as $r\to\infty$.
@@ VIII Discussion
Heuristically, the reduction of local entanglement can be attributed to a kind of entanglement monogamy. Local modes are more entangled with their partners, but the partner modes are not accessible to cosmological observers. The strong entanglement with the partner mode prevents the local mode from becoming substantially entangled with other local modes.
@@ Acknowledgements.
The content of this paper has benefited from discussions with A. Ashtekar, E.Bianchi, A. Delhom, J. Martin, E. Martin-Martinez, S. Nadal, W. van Suijlekom, V. Vennin and K. Yamaguchi. PRM thanks the Perimeter Institute for Theoretical Physics for their support as well as LSU for hosting her for a visit during which part of this work was done. PRM also thanks the Julian Schwinger Foundation for financial support at the 2025 Peyresq Spacetime Meeting, where useful discussions influenced this work. P.R.M. is supported by the 1851 Research Fellowship. This research has been partially supported by the Blaumann Foundation. I.A. is supported by the NSF grants PHY-2409402 and PHY-2110273, by the RCS program of Louisiana Boards of Regents through the grant LEQSF(2023-25)-RD-A- 04, by the Hearne Institute for Theoretical Physics and by Perimeter Institute of Theoretical Physics through the Visitor fellow program. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. This article was finished while BB was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).

### v2

@@ I Introduction
Beyond its implications for cosmology, this result emphasizes the need to distinguish sharply between entanglement and correlation in quantum field theory. The intuitive link between the two often stems from our experience with pure states, where all correlations between a subsystem and its complement are necessarily due to entanglement. In field theory, however, the situation is more subtle. In particular, the reduced state of any field mode that is localized in space—i.e., has compact spatial support—is necessarily mixed (this is a direct consequence of the Reeh-Schlieder theorem [81]). For mixed states, there is no straightforward relationship between correlation and entanglement.
The entanglement content of de Sitter space has been extensively explored in the literature [36, 5, 55, 52, 53, 54, 78, 61, 9, 67, 21, 20, 18, 19, 14, 60, 25, 57, 58, 61, 72, 62, 35, 16, 17, 59, 3, 67, 29, 13, 67, 64, 59, 51, 71, 69, 70, 11], raising the natural question of what new insights this article seeks to provide. Previous studies have approached the problem from diverse perspectives and using a variety of tools, each with its own strengths and limitations—and sometimes producing results that appear in tension with one another. One prominent approach involves quantifying the entanglement between a region and its complement using entanglement entropy [57]. This method has led to interesting findings, suggesting that the entire de Sitter space contains more entanglement than Minkowski space—a result we independently confirm using a different methodology.
@@
The analysis presented here uncovers structural aspects of entanglement in the cosmological patch of de Sitter spacetime that, to our knowledge, have not been previously explored—although some elements of our conclusions were anticipated in [70], based on the study of a restricted family of non-compactly supported field modes. In particular, by employing the concept of partner modes [48, 94, 39, 1], we are able to characterize the subtle manner in which entanglement is distributed in the Bunch-Davies vacuum, revealing important differences from Minkowski spacetime. This, in turn, clarifies that the apparent tension between results from entanglement entropy and entanglement harvesting protocols is not a contradiction. On the contrary, we demonstrate that these findings are closely connected, and in fact, one is a consequence of the other.
For a summary of some of the results presented here, see [83]. The rest of the paper is organized as follows. In Sec. II, we summarize the tools that will be used in the rest of the paper. Sec. III applies these tools to the symmetric vacuum in the cosmological patch of de Sitter spacetime, and discusses some important features of correlation functions together with a few illustrative examples. In Sec. IV, we analyze the von Neumann entropy of a single field degree of freedom. Sections V and VI analyze mutual information and entanglement, respectively, between pairs of compactly supported modes of the field. Finally, in Sec. VII we focus on the notion of partners: the system that codifies the correlations between a local mode and the rest of the field. We use this notion to uncover the distribution of entanglement between different local modes of a scalar field in the Bunch-Davies vacuum in Sec. VII.2. Finally, Sec. VIII summarizes our findings and discusses their implications for curved spacetime quantum field theory and observational signatures of entanglement in the early universe. Throughout this paper, we use units in which $\hbar=c=1$ and adopt the following convention for the Fourier transform $\tilde{h}(\vec{k})=\int d^{3}x\,h(\vec{x})e^{i\vec{k}\cdot\vec{x}}$.
@@ III The cosmological patch of de Sitter spacetime
The de Sitter group in four spacetime dimensions has ten independent Killing vector fields. Locally, all of them are isometries of PdS. However, because the Poincaré patch covers only a portion of de Sitter space, not all of these transformations correspond to global isometries of PdS. Only the subgroup that leaves the Poincaré patch invariant corresponds to global isometries of PdS (see, e.g., [7, Sec. IV C]). This subgroup forms a seven-dimensional Lie group, generated by three spatial translations, three rotations, and one additional isometry defined by the Killing vector field
@@
In the flat-space limit $H\to 0$ and $m\to 0$ (in this order), the terms $\mathcal{N}_{\Phi\Phi}$, $\mathcal{N}_{\Pi\Pi}$, and $\mathcal{N}_{\Phi\Pi}$ vanish, and (45)-(47) reduce to the smeared two-point functions of Minkowski spacetime, as given in Eq. (II.2.1).
@@
It is also worth noting that condition (50) is morally analogous to introducing an infrared cutoff $k_{c}$ on the modulus of the wave numbers in Fourier space. However, (50) is a somewhat milder constraint, as it corresponds to removing only the zero mode $\vec{k}=0$, rather than truncating all low-momentum modes $|\vec{k}|<k_{c}$. [^5] A subset of these special modes has been recently examined in [70] within the framework of a massless scalar field in de Sitter spacetime. This remark serves to contextualize the rather drastic limitation on accessible degrees of freedom imposed by introducing an infrared cutoff.
@@ IV Von Neumann entropy
Case 1: Using Eqs. (45)- (47), we find $\nu_{A}^{2}-(\nu_{A}^{\mathrm{Mink}})^{2}=\mathfrak{a}\,(RH)^{4-4\mu^{2}}(1+\mathcal{O}(\mu^{2}))+\mathcal{O}[(RH)^{3-3\mu^{2}}]\,$, with
@@
Since homogeneous Sobolev spaces of order $|s|<3/2$ are Hilbert spaces with an inner product given in Eq. (92[10], the Cauchy-Schwarz inequality implies that $\mathfrak{a}\geq 0$. Saturation of this inequality, i.e., $\mathfrak{a}=0$, can only occur if $f_{A}^{(1)}=f_{A}^{(2)}$, which corresponds to Case 3, and will be analyzed below. Thus, $\nu_{A}^{2}-(\nu_{A}^{\mathrm{Mink}})^{2}>0$ for $RH\gg 1$, $mH\ll 1$ whenever $0\neq f^{(1)}_{A}\neq f^{(2)}_{A}\neq 0$. Furthermore, since $\mathfrak{a}>0$, $\nu_{A}^{2}-(\nu_{A}^{\mathrm{Mink}})^{2}$ —and hence $\nu_{A}$—grows with $H$, which automatically implies that the entropy of the mode $A$ increases as $H$ increases.
@@ V.1 Examples
Figure 7: Geometric configuration of the modes in Example 6. The shell has inner and outer radii $R_{S}\pm d$, where $2d$ is its width and $R_{S}=R+\Delta x$. Here, $\Delta x$ denotes the radial distance between the edge of the ball (of radius $R$), and the inner edge of the shell.
@@ VI.1 Examples
This case is interesting because $\tilde{\nu}_{-}^{\mathrm{Mink}}<1$ for a certain range of $\Delta x$ (see Fig. 10), meaning that the two modes are entangled in the Minkowski vacuum. In this ball-shell configuration, the two modes are “closer together” compared to the ball-ball configuration, since the shell completely surrounds the ball, which intuitively explains why there is more entanglement.
@@
    LN decays exponentially with the radial distance $\Delta x$ between the shell and ball modes, vanishing beyond a certain threshold distance;
@@ VII Partner systems and entanglement distribution
The apparent contradiction is resolved by analyzing how correlations and entanglement are spatially distributed in the Bunch–Davies vacuum. In particular, we show that the notion of a partner mode [48, 94, 39, 1] provides a convenient framework to characterize the spatial structure of entanglement in quantum field theory in the cosmological patch of de Sitter spacetime. Partner modes were previously discussed in de Sitter spacetime in [70] for a subset of special modes belonging to the family introduced in a Remark on page III. Here, we discuss universal aspects of partner modes in the cosmological patch of de Sitter spacetime.
@@ VII.1 Asymptotic behavior of the partner of a local mode in de Bunch–Davies vacuum
More precisely, let $\gamma_{A_{p}}$ be any vector in $\Gamma_{A_{p}}^{\mathbb{C}}$—the complex version of $\Gamma_{A_{p}}$—with unit Klein–Gordon norm (see footnote 7), so that the pair $(\gamma_{A_{p}},\gamma_{A_{p}}^{*})$ forms a symplectically-orthonormal basis of $\Gamma^{\mathbb{C}}_{A_{p}}$, from which we can obtain a basis in $\Gamma_{A_{p}}$. If we write $\gamma_{A_{p}}(\vec{x})=(g_{p}(\vec{x}),f_{p}(\vec{x}))$, then $g_{p}(\vec{x})$ falls off at least as $r^{-2\mu^{2}}$ and $f_{p}(\vec{x})$ falls off as $r^{-1-\mu^{2}}$, as $r\to\infty$.
@@ VIII Discussion
Heuristically, the reduction of local entanglement can be attributed to a kind of entanglement monogamy. Local modes are more entangled with their partners, but the partner modes are not accessible to cosmological observers. The strong entanglement with the partner mode prevents the local mode from becoming substantially entangled with other local modes. The link between entanglement monogamy and the local thermal properties of the Bunch Davies vacuum has been pointed out in [70] based on the analysis of a rather special family of non-compactly supported field modes (see the Remark on page III for the definition of the family to which these modes belong to).
@@
Our analysis relies on the Gaussian nature of the Bunch–Davies vacuum, which is widely regarded as an excellent approximation to the quantum state of cosmological perturbations. It is therefore natural to ask how our conclusions might be affected by the presence of non-Gaussian features. In general, non-Gaussian states can exhibit entanglement properties that differ qualitatively from those of Gaussian states, and in principle could lead to an enhancement, or even more degradement, of entanglement between local modes. However, observational constraints on primordial non-Gaussianity from the cosmic microwave background and large-scale structure indicate that any such effects must be very small [4]. We therefore expect that non-Gaussian corrections compatible with current observational bounds would not qualitatively modify the results presented here. A detailed quantitative analysis of genuinely non-Gaussian states, however, lies beyond the scope of the present work.
@@ Acknowledgements.
The content of this paper has benefited from discussions with A. Ashtekar, E.Bianchi, A. Delhom, J. Martin, E. Martin-Martinez, S. Nadal, W. van Suijlekom, V. Vennin and K. Yamaguchi. P.R.-M. thanks the Perimeter Institute for Theoretical Physics for their support as well as LSU for hosting her for a visit during which part of this work was done. P.R.-M. also thanks the Julian Schwinger Foundation for financial support at the 2025 Peyresq Spacetime Meeting, where useful discussions influenced this work. P.R.-M. is supported by the 1851 Research Fellowship. This research has been partially supported by the Blaumann Foundation. I.A. is supported by the NSF grants PHY-2409402 and PHY-2110273, by the RCS program of Louisiana Boards of Regents through the grant LEQSF(2023-25)-RD-A- 04, by the Hearne Institute for Theoretical Physics and by Perimeter Institute of Theoretical Physics through the Visitor fellow program. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. This article was finished while B.B. was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).